A stock has an expected return of 12 percent, its beta is 0.40, and the risk-free rate is 7.2 percent. The expected return on the market must be

options 19.2% 20.7% 16.3% 18.9% 22.4%


A stock has a beta of 1.55, the expected return on the market is 16 percent, and the risk-free rate is 9.5 percent. The expected return on this stock must be

options: 21.6% 19.6% 34.8% 28.6% 24.5%
Answer #1

expected return=risk free rate+beta*(market rate-risk free rate)

a.

12=7.2+0.4*(market rate-7.2)

(12-7.2)=0.4*(market rate-7.2)

market rate=(12-7.2)/0.4+7.2

=19.2%

b.Expected rate=9.5+1.55*(16-9.5)

=19.6%(Approx).